The perimeter of a rectangle is 114 centimeters. If the ratio of its width to its lenght is 7 to 12, find the dimensions of the rectangle.
Answer by checkley77(12844)

P=2L+2W
114=2*7N+2*12N N IS THE MULTIPLIER FOR THE RATIO VALUES.
114=14N+24N
114=38N
38N=114
N=114/38
N=3 ANS. FOR THE MULTIPLIER.
7*3=21 FOR THE LENGTH.
12*3=36 FOR THE WIDTH.
PROOF:
114=2*21+2*36
114=42+72
114=114